OpenStudy (anonymous):

yo! when u wanna determine your x and y intercept for a graph. and u 1st set y=0 and got -1 for x do u also set x=0 to find your y? or =-1?

OpenStudy (anonymous):

set x=0 to find your y-intercept. set y=0 to find your x-intercept.

OpenStudy (anonymous):

look at this funtion f(x)=2(x+1)/x^2 if i do dat,my y intercept turns out to be infinity

OpenStudy (anonymous):

am confuse,am doing calculus,trying to plot a graph

OpenStudy (anonymous):

anytime you want to find x and y intercepts that's what you would do. but you also have to consider the domain of your function first. notice that the function you listed is a rational function. these type of functions need special attention.

OpenStudy (y2o2):

0 doesn't belong to the domain of this function so it doesn't have a y intercept.

OpenStudy (anonymous):

my domain is 0

OpenStudy (anonymous):

you know dividing by zero is not allowed. for example 5/0 = ? so in your rational function, you must not let the denominator be zero. so x cannot be zero in the first place.

OpenStudy (anonymous):

no, your domain is any number except 0.

OpenStudy (anonymous):

like you said you were trying to graph.. you got values f(1), f(2), f(2), f(-1), f(-2)... but what happened when you tried f(0)?

OpenStudy (anonymous):

yeah sorry lol,(x not =0) u are right so took my limits in the negative and positive.my vertical asymptotes ended up as positive and my horizontal asympto =0

OpenStudy (anonymous):

f(0) gave me =2/0

OpenStudy (anonymous):

right, 2/0 has no meaning. so you can't graph the point (0, f(0)).

OpenStudy (anonymous):

so what happens now am doing graph,of application of Differentiation

OpenStudy (anonymous):

whatever x value makes your denominator zero will be your vertical asymptote (MOST of the time)

OpenStudy (anonymous):

try put in values close to x=0 like .5, then .1, then .01, .001. do you get y-values when you stick in those numbers?

OpenStudy (anonymous):

mine is positive infinity in both left and right hand side

OpenStudy (anonymous):

yes, so your graph should look like this: